04 Categorical Variables: Problem 16

 Complete the table below 

Books read within the past yearFrequencyRelative FrequencyCumulative Frequency
none6

1-5
0.216216216216216
6-1016

11-15
0.162162162162162
16-2014

21-2510

Total
1

Note: You can ean partial credit on this problem. 
Answer #1

Let N shows the total of frequency. The frequencies for interval 1-5 and 11-15 are 0.216216216216216N and 0.162162162162162N respectively.

Now,

6+ 0.216216216216216N +16+0.162162162162162N+14+10 = N

0.621621621621622N = 46

N = 46 / 0.621621621621622 = 73.999999999999954956521739130462

That in N is 74.

Now frequency for interval 1-5 is

0.216216216216216 *74 = 15.999999999999984

That is frequency of the interval 1-5 is 16.

Now frequency for interval 11-15 is

0.162162162162162*74 = 11.999999999999988

That is frequency of the interval 11-15 is 12.

Following is the completed table:

Relative frequency, f/74 Books read within the past year none 1-5 6-10 11-15 16-20 21-25 Total Frequency, f Cumulative frequency 6 16 16 12 14 10 74 0.081081081081081 0.216216216216216 0.216216216216216 0.162162162162162 0.189189189189189 0.135135135135135 6 38 50 64